# Zhang, Liu, and Zhao

### v3

@@ 3.1 Construction of Surrogation Functions
      $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies Conditions (P2) and (P3) for every $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ with $\varepsilon\geq 0$;
    2. (b)
      $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies Condition (P1) for every $I\in\mathcal{I}^{0}(\theta^{\prime})$;
    3. (c)
@@
      $\widehat{F}(\cdot;\cdot,\xi,I)$ is continuous on $\Theta\times\Theta$ given a fixed index mapping $I$;
    4. (d)
@@
      the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,I)$ satisfies Condition (P4) for any $\varepsilon\geq 0$.
  3. 2.
    The minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)$ satisfies Condition (P5) for any $\varepsilon\geq 0$ and $\xi\in S$.
To illustrate the specific construction of surrogation functions satisfying the above assumptions, we consider the following three cases of the outer function $\varphi$:
@@
    $\varphi(\cdot;\xi)$ is piecewise affine for any $\xi\in S$.
@@
    $\varphi(\cdot;\xi)$ can be decomposed into monotonic convex functions for any $\xi\in S$.
@@
    $\varphi(\cdot;\xi)$ is smooth with Lipschitz gradient for any $\xi\in S$.
@@
In Appendix 9, we provide the surrogation construction of $F(\cdot;\xi)$ for these cases with verification of Assumptions 1 and 2. The three cases cover a broad class of cost functions. In particular, Case 2 is a multivariate extension of the one considered in Cui et al. (2018), which shows that any univariate convex function can be written as the sum of monotonic convex functions. The surrogation construction for Case 3 is new and suitable for nonconvex smooth cost functions.
3.2 Enhanced SMM Algorithm
We present the enhanced SMM algorithm in Algorithm 1.
@@ 3.2 Enhanced SMM Algorithm
2       for $s=1,\dots,N_{\nu}$, draw $\xi_{\nu}^{s}$ independently and uniformly from $S$;
3       select $I_{\nu}^{\varepsilon}$ uniformly from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$;
4       compute $\theta^{\nu+\frac{1}{2}}\in\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}_{N_{\nu}}(\theta;\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\left\lVert\theta-\theta^{\nu}\right\rVert^{2}\right\}$;
5       update the iterate $\theta^{\nu+1}=\begin{cases}\theta^{\nu+\frac{1}{2}},&\text{if}\ \widehat{F}_{N_{\nu}}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{N_{\nu}}(\theta^{\nu}),\\ \theta^{\nu},&\text{otherwise.}\end{cases}$
@@
Algorithm 1 Enhanced SMM Algorithm
At iteration $\nu$, with samples $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ and an $\varepsilon$-active index mapping $I^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$ randomly selected in Steps 2 and 3, we compute in Step 4 the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ with respect to the upper surrogation function $\widehat{F}_{N_{\nu}}(\cdot;\theta^{\nu},I_{\nu}^{\varepsilon})\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}\widehat{F}(\cdot;\theta^{\nu},\xi_{\nu}^{s},I^{\varepsilon}_{\nu})$ of the sample average approximation function $F_{N_{\nu}}(\cdot)\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}F(\cdot;\xi_{\nu}^{s})$. Since the surrogate function value $\widehat{F}_{N_{\nu}}(\theta^{\nu};\theta^{\nu},I_{\nu}^{\varepsilon})$ could be strictly larger than $F_{N_{\nu}}(\theta^{\nu})$, the candidate point $\theta^{\nu+\frac{1}{2}}$ may not lead to a valid update. Hence, it is accepted in Step 5 only when sufficient descent is satisfied. In the special case of $\varepsilon=0$, the touching condition (P1) is satisfied. Thus the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ naturally satisfies the sufficient descent inequality in Step 5, and we directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$.
It is worth noticing that the randomized index-mapping selection in Step 3 is vital for the ERM problem, without which the number of subproblems to be solved at each iteration, identical with the cardinality of $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, would grow exponentially with the sample size. Furthermore, by combining the enhanced $\varepsilon$-active surrogation with incremental sample averaging, the enhanced SMM can achieve satisfactory descent progress in early iterations with relatively small sample sizes, thus improving the computational efficiency; yet such a combination raises the major challenges of convergence analysis to be addressed in the next section.
4 Convergence Analysis of Enhanced SMM
In this section, we first introduce preliminaries on stationarity concepts and then analyze the convergence properties of the enhanced SMM algorithm for solving problem (9). Since much of the convergence analysis is quite technical, we present the main theoretical results with essential discussions while leaving analytical details in Appendix 10.
4.1 Preliminaries on Stationarity
Suppose that the outer function $\varphi$ is Bouligand-differentiable (B-differentiable), i.e., it is locally Lipschitz continuous and directionally differentiable on $\mathbb{R}^{d}$. Then the composite function $F$ is also B-differentiable according to Cui and Pang (2021, Proposition 4.1.2). We say that $\theta^{d}\in\Theta$ is a d-stationary point of problem (9) if $F^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for any $\theta\in\Theta$. For a general nonconvex problem on a convex set, d-stationarity is identified as the sharpest among the first-order stationarity concepts, which gives a necessary condition for local optimality. Specialized to the composite DC program (9), it is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2 (Cui et al. 2018, Proposition 2), and more generally we further define the following strengthened d-stationarity.
Definition 4.1
Given $\varepsilon\geq 0$ and $\rho>0$, we say that $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (9) if ${\{\theta^{d}\}=\mathop{\arg\min}_{\theta\in\Theta}\left\{\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}\widehat{F}(\theta;\theta^{d},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}}$, where $\widehat{F}(\theta;\theta^{d},I)\coloneqq\mathbb{E}\,[\widehat{F}(\theta;\theta^{d},\widetilde{\xi}_{n},I)\,]$, and $\{\widehat{F}(\cdot;\theta^{d},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}$ satisfy Assumption 1 for every $\xi\in S$.
This concept is generalized from the $(\alpha,\eta)$-d-stationarity defined in Lu et al. (2019) for structured DC programs. Due to the requirement of the unique proximal mapping point in the defining condition, our concept is slightly stronger for a reduced composite DC program with the identity outer function $\varphi$. Moreover, it is weaker than the composite $\varepsilon$-strong d-stationarity introduced in Qi et al. (2021) (see also Cui and Pang (2021, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition that is necessary for the follow-up convergence analysis.
Let $\mathcal{S}^{d}$, $\mathcal{S}^{d}_{\varepsilon,\rho}$, and $\mathcal{S}^{*}$ denote the sets of d-stationary points, composite $(\varepsilon,\rho)$-strongly d-stationary points, and global minimizers of (9), respectively. We show below that the composite strong d-stationarity is stronger than d-stationarity and necessary for global optimality.
@@ 4.1 Preliminaries on Stationarity
Under Assumption 1 for the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ for each $\xi\in S$, it holds for problem (9) that $\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$ for any $\varepsilon\geq 0$ and $\rho>0$. In particular, $\mathcal{S}^{d}_{0,\rho}=\mathcal{S}^{d}$ for any $\rho>0$.
The proof is provided in Appendix 10.1. Furthermore, we can show that $\mathcal{S}^{*}=\mathcal{S}^{d}_{\varepsilon,\rho}$ for sufficiently large $\varepsilon$ and suitable $\rho$ under Assumption 2 and a regularity condition that d-stationary values of problem (9) are isolated from the optimal value, which naturally hold when $\varphi$ is in Case 1 or Case 2; this is formally stated as Proposition 10.2 in Appendix 10.1.
4.2 Convergence Analysis of Enhanced SMM Algorithm
Now we establish the convergence analysis of Algorithm 1 in terms of (composite strong) d-stationarity for problem (9). Throughout this section, we make several blanket assumptions that the outer function $\varphi$ is B-differentiable, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ satisfies Assumption 1 for each $\xi\in S$.
Since the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{d})\coloneqq\underset{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}{\min}\widehat{F}(\theta;\theta^{d},I)$ is only of theoretical interest in defining composite strong d-stationarity, we need an auxiliary residual as a bridge to connect with the sampling-based surrogation function $\widehat{F}_{N_{\nu}}(\cdot;\theta^{\nu},I_{\nu}^{\varepsilon})$ utilized in Algorithm 1 for convergence analysis. With $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ and the corresponding surrogate $\widehat{F}(\cdot;\theta^{\prime},I)$ at $\theta^{\prime}\in\Theta$, we denote the proximal map $\mathrm{prox}_{\widehat{F}/\rho}(\theta^{\prime},I)\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\theta^{\prime},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}$ and further define
$P^{\rho}_{I}(\theta^{\prime})\coloneqq\begin{cases}\mathrm{prox}_{\widehat{F}/\rho}(\theta^{\prime},I)&\text{if}\ \min_{\theta\in\Theta}\,\left\{\widehat{F}(\theta;\theta^{\prime},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}\leq F(\theta^{\prime}),\\ \theta^{\prime}&\text{otherwise}.\end{cases}$ (10)
The map $P^{\rho}_{I}(\theta^{\prime})$ can be regarded as an iterative update of Algorithm 1 accounting for all $n$ scenarios, with which we define the averaged residual
$\bar{r}^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\frac{1}{\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert}\sum_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\left\lVert\theta^{\prime}-P^{\rho}_{I}(\theta^{\prime})\right\rVert.$
The following proposition gives convergence rates of Algorithm 1 in terms of $\bar{r}^{\varepsilon,\rho}$ in expectation with the proof provided in Appendix 10.5.
Proposition 4.3
Let $\widetilde{\theta}^{\,T}$ be the output of Algorithm 1 after $T$ iterations with $\varepsilon\geq 0$, $\eta\geq 0$, and $\{N_{\nu}\}_{\nu}$. Then there exists $C_{1}>0$ such that for any $\rho>\eta$,
@@ 4.2 Convergence Analysis of Enhanced SMM Algorithm
By relating $\bar{r}^{\varepsilon,\rho}$ to another residual that measures the proximal mapping gap with respect to $\widehat{F}_{\min}^{\varepsilon}(\cdot;\theta^{\prime})$, we can establish the convergence of Algorithm 1 to composite strong d-stationarity. The error-bound analysis for the nonasymptotic convergence to d-stationarity needs the following assumption generalized from Drusvyatskiy and Lewis (2018). We denote the closure operator of a set by $\mathrm{cl}(\cdot)$ and $\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)\coloneqq\mathop{\arg\min}_{\theta^{\prime}\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta^{\prime};\theta)+\frac{\rho}{2}\left\lVert\theta^{\prime}-\theta\right\rVert^{2}\right\}$.
  1. [label=(B0), resume*=B]
  2. 1.
The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Appendix 10.4. It can be expected to hold in more general cases since it is a local property only requiring the convexity and quadratic growth of $F$ near $\mathcal{S}^{d}$.
Theorem 4.4 (Convergence of Enhanced SMM)
For problem (9), let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon>0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. Then the following results hold:
  1. (a)
@@
    Any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of problem (9) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$.
  2. (b)
    Under Assumptions 2 and 1, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$. In particular, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, the result holds without the two assumptions.

### v4

@@ 3.2 Enhanced sampling-based MM algorithm for learning PADR
    $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ satisfies (P2) and (P3) for every $I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)$ with $\varepsilon\geq 0$;
  3. (B3)
@@
    $\widehat{F}(\cdot,\xi;\,\cdot,I)$ is continuous on $\Theta\times\Theta$ given fixed indices $I=(I_{1},I_{2})\in[K]^{2}$;
  4. (B4)
@@
    the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}\widehat{F}(\theta,\xi;\bar{\theta},I)$ satisfies the directional derivative consistency for any $\varepsilon\geq 0$; namely, $\widehat{F}_{\min}^{\varepsilon}(\cdot\,;\bar{\theta})$ is directionally differentiable on $\Theta$ and satisfies $\widehat{F}_{\min}^{\varepsilon}(\cdot\,;\bar{\theta})^{\prime}(\bar{\theta};d)=F^{\prime}(\bar{\theta};d)$ for any $d\in\mathbb{R}^{q}$.
The directional derivative consistency condition is introduced in Cui and Pang (2021) to relate the directional derivatives of surrogation and original functions. We consider the following three cases of the outer function $\varphi$ covering a broad class of cost functions.
@@
    $\varphi(\cdot,\xi)$ is piecewise affine for any $\xi\in\Xi$.
@@
    $\varphi(\cdot,\xi)$ can be decomposed into monotonic convex functions for any $\xi\in\Xi$.
@@
    $\varphi(\cdot,\xi)$ is smooth with Lipschitz gradient for any $\xi\in\Xi$.
@@
For the above cases, we can construct the surrogate function of $F(\cdot,\xi)$ satisfying Assumptions (B1) to (B4), by utilizing surrogate functions (12) and (13) of the inner PA function $f(\cdot,\xi)$ with the selected index $I=(I_{1},I_{2})$. The specific construction and verification are provided in Online Appendix 7. In particular, Case 2 is a multivariate extension of the one considered in Cui et al. (2018). Case 3 is suitable for nonconvex smooth cost functions that are applicable for many practical problems.
Proposition 3.2
Let $\Theta$ be a compact set. For Case 1 to Case 3, given $\varepsilon\geq 0$ and $\bar{\theta}\in\Theta$, there exists a family of enhanced surrogate functions based on the index set $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$ of the inner PA function that satisfy Assumptions (B1)-(B4) for any $\xi\in\Xi$.
3.2.2 Concepts on stationarity.
For a general nonconvex problem on a convex set with the directional differentiable objective, d-stationarity is identified as the sharpest kind among the first-order stationarity concepts. Specialized to the composite DC program (10), according to Cui et al. (2018, Proposition 2), d-stationarity is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2. Lu et al. (2019) proposes the strengthened d-stationarity for structured DC programs which takes into account the $\epsilon$-active index set. We extend to the concept of composite $(\varepsilon,\rho)$-strong d-stationarity specialized for the PADR-based ERM problem (9). The composite strongly d-stationary point can be interpreted as the fixed point corresponding to the proximal mapping of the expected minimum surrogate function $\mathbb{E}[\,\widehat{F}^{\varepsilon}_{\min}(\theta,\widetilde{\xi}^{n};\bar{\theta})]$ based on the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfying Assumptions (B1)-(B4).
Definition 3.3
Given $\varepsilon\geq 0$ and $\rho>0$, we say that $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (10) if ${\{\theta^{d}\}=\mathop{\arg\min}_{\theta\in\Theta}\left\{\mathbb{E}[\,\widehat{F}^{\varepsilon}_{\min}(\theta,\widetilde{\xi}^{n};{\theta}^{d})\,]+\frac{\rho}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}}$. Correspondingly, let $\mathcal{S}^{d}_{\varepsilon,\rho}$ denote the set of composite $(\varepsilon,\rho)$-strongly d-stationary points.
By Condition (B4), we must have $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}_{d}$ for any $\varepsilon\geq 0$ where $\mathcal{S}_{d}$ denotes the set of d-stationary points. Moreover, with $\varepsilon>0$, we have the inclusion property $\mathcal{I}^{0}(x,\xi)\subseteq\mathcal{I}^{0}(\hat{x},\xi)\subseteq\mathcal{I}^{\varepsilon}(x,\xi)$ for $x$ sufficiently close to $\hat{x}$, which is crucial for the convergence analysis of the limit point in terms of (composite strong) d-stationarity. With $\mathcal{S}^{*}$ being the set of global minima of (10), the following proposition gives the relation among $\mathcal{S}^{*},\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{d}$. The proof is provided in Online Appendix 8.1.
Proposition 3.4
$\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$ for any $\varepsilon\geq 0$ and $\rho>0$. In particular, $\mathcal{S}^{d}_{0,\rho}=\mathcal{S}^{d}$ for any $\rho>0$.
Due to the requirement of a unique proximal mapping point in the defining condition, our concept is slightly stronger than $(\alpha,\eta)$-d-stationarity defined in Lu et al. (2019) when the problem is reduced to composite DC program. Moreover, it is weaker than composite $\varepsilon$-strong d-stationarity introduced in Qi et al. (2021b) (see also Cui and Pang (2021, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition which turns out to be essential for the algorithm development and the follow-up convergence analysis. Furthermore, in Online Appendix 8.1, we show that $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ under a regularity assumption on $F$. It implies that the definition of $\mathcal{S}^{d}_{\varepsilon,\rho}$ serves as a natural connection between d-stationarity and global optimality.
3.2.3 ESMM algorithm and convergence analysis.
We propose the ESMM algorithm with a natural idea of combining the randomized $\varepsilon$-active surrogation with sequential sampling presented in Algorithm 1. At iteration $\nu$, a subset of samples $\Xi_{\nu}=\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ of size $N_{\nu}$ is generated in step 2. Then in step 3 a sampling-based surrogate function $\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ is constructed which can be regarded as the upper surrogation function to the sample average approximation function $F_{\nu}(\cdot)\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}F(\cdot,\widetilde{\xi}_{\nu}^{s})$ with the combination of indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}=\{\widetilde{I}^{\varepsilon}_{\nu,s}\}_{s=1}^{N_{\nu}}$ randomly selected from the $\varepsilon$-active index sets $\{\mathcal{I}^{\varepsilon}(\theta^{\nu},\widetilde{\xi}_{\nu}^{s})\}_{s=1}^{N_{\nu}}$, respectively. The randomized selection combined with the sampling scheme enables the algorithm to solve subproblems of small sizes in the early iterations. By the convexity of the surrogate function, in step 4 the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ can be obtained via the state-of-the-art convex programming solvers. However, since the indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}$ are randomly selected from the $\varepsilon$-active index sets, the touching condition (P1) may not be satisfied when $\varepsilon>0$; namely, the surrogate function value $\widehat{F}_{\nu}(\theta^{\nu};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ might be strictly larger than $F_{\nu}(\theta^{\nu})$. Thus the candidate solution $\theta^{\nu+\frac{1}{2}}$ may not lead to a valid update with the desired descent property. Hence, in step 5 the candidate solution $\theta^{\nu+\frac{1}{2}}$ is accepted only when sufficient descent is satisfied. In the special case of $\varepsilon=0$ for which the touching condition (P1) is satisfied, the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ satisfies the sufficient descent inequality in step 5, and thus we could directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$.
@@
2       generate $\Xi_{\nu}\coloneqq\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ independently and uniformly from $\Xi$;
3       construct the surrogate function $\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}\,\widehat{F}(\theta,\widetilde{\xi}_{\nu}^{s};\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu,s})$ where $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}\coloneqq\{\widetilde{I}^{\varepsilon}_{\nu,s}\}_{s=1}^{N_{\nu}}$ and $\widetilde{I}^{\varepsilon}_{\nu,s}$ is uniformly generated from $\mathcal{I}^{\varepsilon}(\theta^{\nu},\widetilde{\xi}_{\nu}^{s})$;
4       compute $\theta^{\nu+\frac{1}{2}}\in\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})+\frac{\eta}{2}\left\lVert\theta-\theta^{\nu}\right\rVert^{2}\right\}$;
5       update the iterate $\theta^{\nu+1}=\begin{cases}\theta^{\nu+\frac{1}{2}},&\text{if}\ \widehat{F}_{\nu}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{\nu}(\theta^{\nu}),\\ \theta^{\nu},&\text{otherwise.}\end{cases}$
@@
Algorithm 1 ESMM algorithm
Next, we establish the convergence analysis of Algorithm 1 in terms of (composite strong) d-stationarity for problem (10). Due to the combination of sampling and enhanced technique, we introduce the averaged residual for convergence analysis, which is technically different from the analysis in Pang et al. (2016), Cui et al. (2018), Lu et al. (2019). We present the main theoretical results with essential discussions while leaving technical details in Online Appendix 8. Throughout the analysis, we make several blanket assumptions: the outer function $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$ and any $\bar{\theta}\in\Theta$.
With $\Xi=\{\xi^{s}\}_{s\in[n]}=\{(x^{s},y^{s})\}_{s\in[n]}$, let $\mathcal{I}^{\varepsilon}(\bar{\theta})$ denote the set of all $\varepsilon$-active index combinations of $\{(g(\bar{\theta},\xi),h(\bar{\theta},\xi)\}_{\xi\in\Xi}$, i.e.,
$\mathcal{I}^{\varepsilon}(\bar{\theta})\coloneqq\Big{\{}{\mathbf{I}}^{\varepsilon}=\{(I_{1,s}^{\varepsilon},I_{2,s}^{\varepsilon})\}_{s=1}^{n}\mid I^{\varepsilon}_{1,s}\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi^{s}),I^{\varepsilon}_{2,s}\in\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi^{s}),\,\forall\,s=1,\ldots,n\Big{\}},$
with $|\mathcal{I}^{\varepsilon}(\bar{\theta})|$ representing the total number of combinations of $\varepsilon$-active index selection for the ERM objective function $\frac{1}{n}\sum_{s=1}^{n}F(\theta,\xi^{s})$. For the purpose of analysis, we introduce the surrogate function for such an ERM function, and adjusted proximal mapping point which is motivated by the iterative update in step 5 of Algorithm 1. At $\bar{\theta}\in\Theta$ with $\mathbf{I}^{\varepsilon}=\{I_{s}^{\varepsilon}\}_{s\in[n]}\in\mathcal{I}^{\varepsilon}(\bar{\theta})$, let
$\begin{array}[]{ll}\widehat{F}(\cdot;\bar{\theta},\mathbf{I}^{\varepsilon})&\coloneqq\displaystyle{\frac{1}{n}}\sum_{s=1}^{n}\widehat{F}(\theta,\xi^{s};\bar{\theta},I^{\varepsilon}_{s})\\[14.45377pt] P^{\rho}_{\mathbf{I}^{\varepsilon}}(\bar{\theta})&\coloneqq\begin{cases}\mathrm{prox}_{\widehat{F}/\rho}(\bar{\theta},\mathbf{I}^{\varepsilon})&\text{if}\ \min_{\theta\in\Theta}\,\left\{\widehat{F}(\theta;\bar{\theta},\mathbf{I}^{\varepsilon})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}\leq F(\bar{\theta}),\\ \bar{\theta}&\text{otherwise}.\end{cases}\end{array}$ (15)
where the proximal map $\mathrm{prox}_{\widehat{F}/\rho}(\bar{\theta},\mathbf{I}^{\varepsilon})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\bar{\theta},\mathbf{I}^{\varepsilon})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}$. With the above prerequisite, we introduce the averaged residual
$\bar{r}^{\varepsilon,\rho}(\bar{\theta})\coloneqq\frac{1}{\left\lvert\mathcal{I}^{\varepsilon}(\bar{\theta})\right\rvert}\sum_{\mathbf{I}^{\varepsilon}\in\mathcal{I}^{\varepsilon}(\bar{\theta})}\left\lVert\bar{\theta}-P^{\rho}_{\mathbf{I}^{\varepsilon}}(\bar{\theta})\right\rVert,$
which plays a central role in the follow-up convergence analysis. Under the sampling size sequence $\{N_{\nu}\}$, we present the convergence result of $\bar{r}^{\varepsilon,\rho}$ in expectation in the proposition below, with the proof provided in Online Appendix 8.5.
@@
Let $\widetilde{\theta}^{\,T}$ be the output of Algorithm 1 after $T$ iterations with $\varepsilon\geq 0$, $\eta\geq 0$, and $\{N_{\nu}\}_{\nu=1}^{T}$. There exists $C_{1}>0$ such that for any $\rho>\eta$,
@@
By relating $\bar{r}^{\varepsilon,\rho}$ to the proximal mapping gap with respect to $\widehat{F}_{\min}^{\varepsilon}(\cdot;\bar{\theta})$, with $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$, we can establish the convergence of Algorithm 1 in terms of composite strong d-stationarity.
Theorem 3.6 (Asymptotic convergence of ESMM)
@@
For problem (10), let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon>0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. Then any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ which must exist is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of problem (10) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$
Theorem 3.6 shows the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity. Moreover, under the aforementioned settings where $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ and a regularity assumption on $F$, any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a global minimizer of (10) with probability 1. This indicates that the PADR-based ERM model together with the ESMM algorithm is asymptotically optimal to the original problem (2). However, evidenced by the numerical implementations, when $\varepsilon$ is large, the candidate solutions produced in step 4 of Algorithm 1 may be rejected by the update rule in step 5 very often, which results in inefficient computation and slow convergence. Such a computational issue is resolved by a shrinking-$\varepsilon$ strategy which will be specified in Section 4.